EXCEL MASTER IT! PROBLEM Using Excel to find the marginal tax rate can be accomplished using the VLOOKUP func- tion. However, calculating the total tax bill is a little more difficult. Below we have shown a copy of the IRS tax table for an unmarried individual for 2018. Often, tax tables are presented in this format. xlsx If taxable income is But not over . $ 9,525 38,700 93,700 195,450 424,950 426,700 The tax is: over... 0 9,526 38,701 93,701 195,451 424,951 426,701 10% of the amount over $0 $952.50 plus 15% of the amount over $9,525 5,328.75 plus 25% of the amount over $38,700 $19,078.75 plus 28% of the amount over $93,700 $47,568.75 plus 33% of the amount over $195,450 $123,303.75 plus 35% of the amount over $424,950 $1 23.9 16.25 plus 39.6% of the amount over $426,700 In reading this table, the marginal tax rate for taxable income less than $9,525 is 10 percent. If the taxable income is between $9.525 and $38,700, the tax bill is $952.50 plus the marginal taxes. The marginal taxes are calculated as the taxable income minus $9.525 times the marginal tax rate of 15 percent.
Chapter 2 Financial Statements, Taxes, and Cash Flow 47 Below, we have the corporate tax table that was applicable 2017 and as shown in Table 2.3. Taxable income is greater than But less than. or equal to... 0 50,001 75,001 100,001 335,001 10,000,001 15,000,001 18,333,334 $50,000 75,000 100,000 335,000 10,000,000 15,000,000 18,333,333 Tax rate 15% 25 34 39 34 35 38 35 a. Create a tax table in Excel for corporate taxes similar to the individual tax table shown above. Your spreadsheet should then calculate the marginal tax rate, the average tax rate, and the tax bill for any level of taxable income input by a user b. For a taxable income of $1,350,000, what is the marginal tax rate? c. For a taxable income of $1,350,000, what is the total tax bill? d. For a taxable income of $1,350,000, what is the average tax rate?
Please answer in excel.
Please show the formulas
Answer #1

Answer:

a)

Tax calculation sample for $100,001-$335,000 slab

Tax=15%*50000+(75000-50000)*25%+(100000-75000)*34% Plus 39% of amount over $100,000.

Tax=$22,250 plus 39% of amount over $100,000.

If taxable income is over But Not over The tax is
$0 $ 50000 15% of amount over $0
50,001 75,000 $7500 plus 25% of amount over $50,000
75,001 100,000 $13750 plus 34% of amount over $75,000
100,001 335,000 $ 22250 plus 39% of amount over $100,000
335,001 10,000,000 $ 113900 plus 34% of amount over $335,000
10,000,001 15,000,000 $340,000 plus 35% of amount over $10,000,000
15,000,001 18,333,333 $2,090,000 plus 38% of amount over $15,000,000
18,333,334 $3,356,666.54 plus 35% of amount over $18,333,333

B) for income $ 1,350,000 the marginal tax rate is 34%.

C) for income $ 1,350,000 the total tax bill is 113900+(1350000-335000)*34%=$459,000

D) for income $ 1,350,000 average tax rate = Total tax/income= 459000/1350000=34%